OpenStudy (anonymous):

6 (-2x+1)= -3(4x+2) last equation I don't understand.

OpenStudy (mathstudent55):

Distribute oin both sides: - 12 x + 6 = - 12x - 6 Add 12 x to both sides: -12x + 6 = -12 x - 6 + 12x +12x 6 = - 6 Since 6 = -6 is false, htere is no solution to this equation

OpenStudy (anonymous):

Ohh. c: thanks.

OpenStudy (mathstudent55):

You're welcome